a car reached raipur from ranchi in 35 minutes with an average speed of 69km/hour.if average speed is increased by 36 km/hr how much time will it take to cover same distance
Answer by ptfile(81)

Let 35min = the duration of the car traveled
69kph = the speed of the car
A car reached raipur from ranchi in 35 minutes with an average speed of 69km/hour.
Distance = (rate)(time)
d = (69)(35)
d = 2415/60
d = 40.25

If average speed is increased by 36 km/hr
40.25 = (69+36)t
40.25 = 105t
(.38)(60)= t
23 = t
The it took to reach raipur with a speed of 105kph is 23 minutes.